OpenStudy (anonymous):

f(x)=log5x Value of f(1/25)?

OpenStudy (anonymous):

Put \(\frac{1}{25}\) where \(x\) is.

OpenStudy (mertsj):

\[f(\frac{1}{25})=\log 5(\frac{1}{25})=\log \frac{1}{5}=\log 5^{-1}\]

OpenStudy (anonymous):

:/ My options are: a. sqrt2 b. -2 c. -1/2 d. 1/2

OpenStudy (mertsj):

Well then, I guess you have posted the problem incorrectly.

OpenStudy (anonymous):

The problem is actually \[f(x)=\log_{5}(x)\] But we are not supposed to solve this for @mathsux4real. What I said should tell him enough to be able to do this problem, unless he is unfamiliar with manipulating exponents.

OpenStudy (mertsj):

I would guess that 5 is the base. But that is not the way you posted it.

OpenStudy (anonymous):

You are right Mertsj. 5 is the base. My fault that I didn't post it that way.

OpenStudy (anonymous):

I meant to type f(x)=log_5x Value of f(1/25)?

OpenStudy (anonymous):

@mathsux4real, what is so hard about this for you?

OpenStudy (anonymous):

Got it. -2.

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

You're welcome!